Jump to main content.


Drake Chemical

EPA Daily Situation Reports for April 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Friday, April 30, 1999

No report available.


Thursday, April 29, 1999

No report available.


Wednesday, April 28, 1999

No report available.


Tuesday, April 27, 1999

No report available.


Monday, April 26, 1999

No report available.


Sunday, April 25, 1999

No report available.


Saturday, April 24, 1999

No report available.


Friday, April 23, 1999

FOR DATES: From 7:00 a.m. Thursday, April 22, 1999 to 7:00 a.m. Friday, April 23, 1999.

WEATHER: Overcast to partly cloudy. The temperature range was 45oF to 69oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, five days a week, operating and maintaining the Thermal Destruction Facility (TDF). All contaminated soils have been remediated. The plant is presently in a hot hold condition awaiting the final approval of last four bottom ash analytical results. The TDF was not on-line yesterday and did not process any material during the time frame of this report. The contractor is performing demobilization related site activities such as ice removal from the oxygen vaporizers, air lancing of internal systems, bag house purging, steam and acid gas scrubber flushing and general maintenance. The waste water treatment plant is operating on a 24 hour schedule until flushing of the scrubber systems have been completed.

II. END OF SITE SOILS FED INTO UNIT: On April 22, 1999, at 6:09 A.M., the feeding of all site soils into the unit was completed.

III. ACTIONS TAKEN: The contractor is performing demobilization related site activities such as ice removal from the oxygen vaporizers, air lancing of internal systems, bag house purging, steam and acid gas scrubber flushing, waste water management and general maintenance. The wastewater treatment plant is operating on a 24-hour schedule until flushing of the steam and acid gas scrubber systems have been completed. Real time perimeter air monitoring (PAM) program is ongoing.

IV. FUTURE PLANS: Major activities for the next week include the completion acid gas scrubber flushing, begin feed preparation liner removal and building demolition, debris handling and offsite shipping, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, inorganic debris decontamination. The dismantling and demobilization of the TDF unit is expected to begin the week of April 26, 1999 with the removal of the oxygen vaporizers and tanks and the taking down of the north flyash bin wall.

V. PROBLEMS NOTED OR ANTICIPATED: None.

VI. NOTED VISITORS: Dr. Scott Tucker

PREPARED BY: John Ricard for Mike Ogden (USACE)


Thursday, April 22, 1999

FOR DATES: From 7:00 a.m. Wednesday, April 21, 1999 to 7:00 a.m. Thursday, April 22, 1999.

WEATHER: Overcast, rain. The temperature range was 32oF to 52oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 22.5 hours and 756.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.6 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, removal of site soils from HDPE liner, equipment decontamination and debris separation.

II. END OF SITE SOILS FED INTO UNIT: On April 22, 1999, at 6:09 A.M., the feeding of all site soils into the unit was completed.

III. ACTIONS TAKEN: The contractor completed feeding of site soils, continued TDF operations, waste water treatment plant operations (WWTP), wastewater management, debris separation, equipment and materials decontamination, rock backfilling, ash handling and ash backfilling. The contractor has completed exposing the HDPE liner, on which the contaminated site soils were staged. Real time perimeter air monitoring (PAM) program is ongoing.

Received bottom ash analytical report (bot-ash-04-17-99-412) for the contents of bin 6. Analytical test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

IV. FUTURE PLANS: Major activities for the next week include the completion acid gas scrubber flushing, begin feed preparation building demolition, debris handling and offsite shipping, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, inorganic debris decontamination. The dismantling and demobilization of the TDF unit is expected to begin the week of April 26, 1999.

V. PROBLEMS NOTED OR ANTICIPATED: None.

VI. NOTED VISITORS: Tom Voltaggio - EPA; Gregg Crystall - EPA; Bill Parsons - PADEP; Mike Welch - PADEP; Larry Newcomer - PADEP; Sandy Downs - PADEP; Lloyd Caldwell - USACE; Denis duBreuil - USACE; Jared Olsen - PADEP; Matt Radek - OHM; Dave Modricker; Dr. Scott Tucker

VII. HAPPY EARTH DAY!

PREPARED BY: John Ricard for Mike Ogden (USACE)


Wednesday, April 21, 1999

FOR DATES: From 7:00 a.m. Tuesday, April 20, 1999 to 7:00 a.m. Wednesday, April 21, 1999.

WEATHER: Overcast, rain. The temperature range was 32oF to 52oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 803.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, removal of site soils from HDPE liner, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant operations (WWTP), wastewater management, feed preparation, debris separation, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor continues to expose and clean the HDPE liner, on which the remaining contaminated sites soils are staged. Real time perimeter air monitoring (PAM) program is ongoing.

Received bottom ash analytical report (bot-ash-04-16-99-411for the contents of bin 4. Analytical test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include the completion of thermal treatment of contaminated soils, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination. The dismantling and demobilization of the TDF unit is expected to begin the week of April 26, 1999.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Larry Newcomer and Sandra Downs (PADEP)

PREPARED BY: Danny Zion for Bill Werntges (USACE)


Tuesday, April 20, 1999

FOR DATES: From 7:00 a.m. Monday, April 19, 1999 to 7:00 a.m. Tuesday, April 20, 1999.

WEATHER: Overcast. The temperature range was 36oF to 53oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 803.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, removal of site soils from HDPE liner, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant operations (WWTP), wastewater management, feed preparation, debris separation, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor continues to expose and clean the HDPE liner, on which the remaining contaminated sites soils are staged. Real time perimeter air monitoring (PAM) program is ongoing.

Received bottom ash analytical reports (bot-ash-04-14-99-409, and bot-ash-04-15-99-410 for the contents of bins 3 and 2. Analytical test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include the completion of thermal treatment of contaminated soils, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination. The dismantling and demobilization of the TDF unit is expected to begin the week of April 26, 1999.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Monday, April 19, 1999

FOR DATES: From 7:00 a.m. Sunday, April 18, 1999 to 7:00 a.m. Monday, April 19, 1999.

WEATHER: Overcast, light rain. The temperature range was 36oF to 52oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 808 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on-line. Off line time was attributable to the daily calibration of the weigh belt. The contractor is performing related site activities such as feed preparation, removal of site soils from HDPE liner, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant operations (WWTP), wastewater management, feed preparation, debris separation, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor continues to expose and clean the HDPE liner, on which the remaining contaminated sites soils are staged. Real time perimeter air monitoring (PAM) program is ongoing.

No new ash analytical results were received during this reporting period. The contractor completed exposing all of the HDPE liner outside of the feed preparations building. All remaining contaminated site soils to be thermally treated are now stockpiled inside of the building.

III. FUTURE PLANS: Major activities for the next week include the completion of thermal treatment of contaminated soils, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination. The dismantling and demobilization of the TDF unit is expected to begin the week of April 26, 1999.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, April 18, 1999

FOR DATES: From 7:00 a.m. Saturday, April 17, 1999 to 7:00 a.m. Sunday, April 18, 1999.

WEATHER: Overcast, light rain. The temperature range was 42oF to 48oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 802.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.6 tons per hour while on-line. Off line time was attributable to the daily calibration of the weigh belt. The contractor is performing related site activities such as feed preparation, removal of site soils from HDPE liner, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant operations (WWTP), wastewater management, feed preparation, debris separation, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor continues to expose and clean the HDPE liner, on which the remaining contaminated sites soils are staged. Real time perimeter air monitoring (PAM) program is ongoing.

Received bottom ash analytical report (bot-ash-04-13-99-408) for the contents of bins 5. Analytical test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received flyash analytical report for (fly-ash-04-13-99) for the contents of bin A. Analytical test results indicated that the thermally treated soils met project requirements for restricted backfill on-site above elevation 553 due to concentrations of selenium, cadmium, and chromium being greater than the DWS but less than 25 times it.

III. FUTURE PLANS: Major activities for the next week include the completion of thermal treatment of contaminated soils, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination. The dismantling and demobilization of the TDF unit is expected to begin the week of April 26, 1999.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Saturday, April 17, 1999

FOR DATES: From 7:00 a.m. Friday, April 16, 1999 to 7:00 a.m. Saturday, April 17, 1999.

WEATHER: Cloudy, overcast with high winds and rain p.m. Friday morning is overcast with light rain. The temperature range was 35oF to 64oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 21.8 hours and 729.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.4 tons per hour while on-line. Off line time was attributable to an unplanned morning Thermal Relief Vent (TRV) opening event caused by a brief interruption in power at 0724 hours and a afternoon thunderstorm event (no TRV). The contractor is performing related site activities such as feed preparation, maintenance on the steam scrubber, system equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant operations (WWTP), wastewater management, feed preparation, debris separation, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor continues to expose and clean the HDPE liner, on which the remaining contaminated sites soils are staged. Real time perimeter air monitoring (PAM) program is ongoing.

Received bottom ash analytical reports (bot-ash-04-11-99-406, and bot-ash-04-12-99-407 for the contents of bins 1and 7. Analytical test results indicated that the thermally treated soils met the project requirements for restricted backfilling on-site above elevation 553 due to the concentration of selenium being greater than the DWS but less than 25 times it.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Thermal treatment of site soils is expected to be complete during the week of April 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: John Ricard for Chad Thompson (USACE)


Friday, April 16, 1999

FOR DATES: From 7:00 a.m. Thursday, April 15, 1999 to 7:00 a.m. Friday, April 16, 1999.

WEATHER: Clear morning, overcast afternoon with light rain Friday morning. The temperature range was 32oF to 64oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.3 hours and 773.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.2 tons per hour while on-line. Off line time was attributable to the contractor performing maintenance on the steam scrubber system. The contractor is performing related site activities such as feed preparation, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor continued to expose and clean the HDPE liner, on which the remaining contaminated sites soils are staged. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-04-08-99-403, and bot-ash-04-10-99-405 for the contents of bins 4 and 2. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

Received bottom ash analytical report (bot-ash-04-09-99-404 for the contents of bins 6. Analytical test results indicated that the thermally treated soil met the project requirements for restricted backfilling on-site above elevation 553 due to the concentration of selenium being greater than the DWS but less than 25 times it. Received fly ash analytical report (fly-ash-4-9-99-080) for the contents of bin D. Analytical test results indicated that the thermally treated soil met the project requirements for restricted backfill on-site above elevation 553 due to concentrations of selenium, cadmium, and chromium being greater than the DWS but less than 25 times it.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Thermal treatment of site soils is expected to be complete during the week of April 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP).

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Thursday, April 15, 1999

FOR DATES: From 7:00 a.m. Wednesday, April 14, 1999 to 7:00 a.m. Thursday, April 15, 1999.

WEATHER: Clear. The temperature range was 32oF to 66oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 6.2 hours and 186 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.2 tons per hour while on-line. Off line time was attributable to the contractor performing monthly interlock checks and equipment calibrations and also due to performing maintenance on the steam scrubber system. The contractor is performing related site activities such as feed preparation, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. The contractor performed monthly interlock checks and calibrations, which included a planned opening of the TRV. The required notifications were made prior to the opening. The contractor used a remote opening device to allow a sample to be collected from the intact stainless steel drum that was unearthed during excavation activities. Maintenance was performed on the steam scrubber unit. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP), Kevin Fanning (CCEMA), and Kurt Davis (TAG Rep.).

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Wednesday, April 14, 1999

FOR DATES: From 7:00 a.m. Tuesday, April 13, 1999 to 7:00 a.m. Wednesday, April 14, 1999.

WEATHER: Clear and sunny. The temperature range was 36oF to 60oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 753.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.5 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, equipment decontamination and debris separation.

Feed to the incinerator was shut down at 0800 hours on April 14, 1999, in order to perform the monthly interlock checks (AWFSOs) and perform the monthly calibrations to the TDF instruments. The interlock checks are being performed to check the TDF's automatic waste feed shutoff systems. As part of the interlock checks, there will be a planned Thermal Relief Vent (TRV) opening scheduled for today. Prior to the TRV opening, waste feed will have been off for several hours and the proper local authorities will be notified.

While the incinerator is shut down for the interlock checks, the opening of the stainless steel drum is to be performed.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment and materials decontamination, decontaminated rock backfilling, ash handling and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program and initiated the OSHA method 93 24-hour air sampling event for BNA at the on-site and off-site air monitoring stations. Roll-off #242, containing PPE and site debris, was shipped off-site as non-hazardous waste to a hazardous waste landfill.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Larry Newcomer and Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for Michael Ogden (USACE)


Tuesday, April 13, 1999

FOR DATES: From 7:00 a.m. Monday, April 12, 1999 to 7:00 a.m. Tuesday, April 13, 1999.

WEATHER: Overcast. The temperature range was 34oF to 56oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 764.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.2 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment and materials decontamination, ash handling and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

Excavation to elevation 545 was completed during this reporting period. All remaining contaminated soils to be thermally treated are staged on top of the HDPE liner which is under the feed preparations and the areas immediately south and east of the building extending to the excavation support system.

Received bottom ash analytical report (bot-ash-04-07-99-402) for the contents of bin 3. Bottom ash analytical reports indicated that the thermally treated soil met the project requirements for restricted backfilling on-site above elevation 553 due to the concentration of selenium being greater than the DWS but less than 25 times it.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. Waste feed to the incinerator will be stopped at 0800 hours on Wednesday April 14, 1999 in order to perform system interlock checks and monthly calibrations. As part of the interlock checks a planned TRV opening will occur during the day on Wednesday.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Chad Thompson for William Werntges (USACE)


Monday, April 12, 1999

FOR DATES: From 7:00 a.m. Sunday, April 11, 1999 to 7:00 a.m. Monday, April 12, 1999.

WEATHER: Overcast, heavy rain at times. The temperature range was 35oF to 44oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 767.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.1 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment and materials decontamination, ash handling and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-04-06-99-401) for the contents of bin 5. Bottom ash analytical reports indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The excavation of the unlined portion of the exclusion zone is scheduled to be complete on Monday, April 12, 1999.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, April 11, 1999

FOR DATES: From 7:00 a.m. Saturday, April 10, 1999 to 7:00 a.m. Sunday, April 11, 1999.

WEATHER: Clear and cool. The temperature range was 34oF to 58oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 767.0 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.1 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, equipment decontamination and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment and materials decontamination, ash handling and ash backfilling. Excavation activities were continued directly south of the CZ113 sheet piling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan. The excavation of the unlined portion of the exclusion zone is scheduled to be complete on Monday, April 12, 1999.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Saturday, April 10, 1999

FOR DATES: From 7:00 a.m. Friday, April 9, 1999 to 7:00 a.m. Saturday, April 10, 1999.

WEATHER: Overcast with rain. The temperature range was 34oF to 45oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 21.9 hours and 689.8 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to the stoppage of feed during a local lightning storm event, as per the site standard operating procedure to prevent power loss to the TDF. The TDF processed contaminated soil at an average rate of 31.5 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment decontamination, rock backfilling, ash handling and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-04-05-99-400A and bot-ash-04-05-99-400B) for the contents of bin 7. Bottom ash analytical reports indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Friday, April 9, 1999

FOR DATES: From 7:00 a.m. Thursday, April 8, 1999 to 7:00 a.m. Friday, April 9, 1999.

WEATHER: Clear and mild. The temperature range was 49oF to 78oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 13.5 hours and 426 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to the stoppage of feed in order to repair several malfunctioning quench tower nozzles. The TDF processed contaminated soil at an average rate of 31.6 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment decontamination, dust control, rock backfilling, ash handling and ash backfilling. Excavation was performed south of the CZ113 sheetpiling. Continued with the real time perimeter air monitoring (PAM) program and completed the OSHA method 93 24-hour air sampling event. The contents of bin 6, which did not meet the organic criteria for backfilling on-site, were returned to the exclusion zone for further thermal treatment.

Received fly ash analytical report (fly-ash-04-04-99-079) for the contents of bin A. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site due to the concentration of cadmium, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, April 8, 1999

FOR DATES: From 7:00 a.m. Wednesday, April 7, 1999 to 7:00 a.m. Thursday, April 8, 1999.

WEATHER: Clear and mild. The temperature range was 45oF to 68oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 759 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.7 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment decontamination, dust control, rock backfilling, shot rock backfilling, ash handling and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program. The contents of bin 4, which did not meet the organic criteria for backfilling on-site, were returned to the exclusion zone for further thermal treatment.

Received bottom ash analytical reports (bot-ash-04-04-99-399A and bot-ash-04-04-99-399B) for the contents of bin 6. The analytical test results indicated that the thermally treated soil stored in bin 6 did not meet the project organic criteria due to the concentration of Beta-Naphthylamine being greater than the organic criteria for backfilling on site. The contents of bin 6 will be returned to the exclusion zone for further thermal treatment.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, April 7, 1999

FOR DATES: From 7:00 a.m. Tuesday, April 6, 1999 to 7:00 a.m. Wednesday, April 7, 1999.

WEATHER: Partly cloudy with periods of rain. The temperature range was 33oF to 60oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 793.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.2 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, lime mixing, equipment decontamination, dust control, rock backfilling, shot rock backfilling, ash handling and ash backfilling. Excavation activities were continued in the exclusion zone south of the CZ113 sheet piling. The contractor continued to expose the HDPE liner on the outside of the feed preparation building on the northern end of the CZ113 sheet piling. Continued with the real time perimeter air monitoring (PAM) program. One roll-off containing the bulked drum contents was shipped off-site and non-hazardous waste to a hazardous waste landfill.

Received bottom ash analytical reports (bot-ash-04-02-99-398A and bot-ash-04-02-99-398B) for the contents of bin 4. The analytical test results indicated that the thermally treated soil stored in bin 4 did not meet the project organic criteria due to the concentration of Beta-Naphthylamine being greater than the organic criteria for backfilling on site. The contents of bin 4 will be returned to the exclusion zone for further thermal treatment.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs, Mike Welch and Larry Newcomer (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, April 6, 1999

FOR DATES: From 7:00 a.m. Monday, April 5, 1999 to 7:00 a.m. Tuesday, April 6, 1999.

WEATHER: Clear and sunny. The temperature range was 30oF to 61oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 765.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.3 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, rock backfilling, ash handling and ash backfilling. Excavation activities were continued in the exclusion zone south of the CZ113 sheet piling. Continued with the real time perimeter air monitoring (PAM) program. Two roll-offs containing PPE, wood debris, metal debris and lime bags were shipped off-site and non-hazardous waste to a hazardous waste landfill.

Received bottom ash analytical report (bot-ash-03-31-99-396) for the contents of bin 1 and reports (bot-ash-4-1-99-397A and bot-ash-4-1-99-397B) for the contents of bin 3. Analytical reports indicated that the thermally treated soils met the project requirements for unrestricted backfilling (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, April 5, 1999

FOR DATES: From 7:00 a.m. Sunday, April 4, 1999 to 7:00 a.m. Monday, April 5, 1999.

WEATHER: Overcast, evening rain. The temperature range was 33oF to 60oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 829.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.7 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, ash handling. Excavation activities were performed in the exclusion zone south of the CZ113 sheet piling. Rock from the screeners was used for backfilling on the 2-foot rock bottom. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-03-28-99-393, bot-ash-03-29-99-394, and bot-ash-03-30-99-395) for the contents of bins 7,5, and 2. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (Fly-ash-3-29-99-078) for the contents of bin C. Analytical test results indicated that the thermally treated soil met the project requirements for backfilling on-site above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for William Werntges (USACE)


Sunday, April 4, 1999

FOR DATES: From 7:00 a.m. Saturday, April 3, 1999 to 7:00 a.m. Sunday, April 4, 1999.

WEATHER: Overcast. The temperature range was 55oF to 74oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 21 hours and 767.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.55 tons per hour while on-line. Down time is contributive to weigh belt calibration and repair work to the 210 and 212 feed hoppers. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, ash handling. Excavation activities were performed in the exclusion zone south of the CZ113 sheet piling. Rock from the screeners was used for backfilling on the 2-foot rock bottom. Also, imported shot rock was placed for the 2-foot rock bottom. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Saturday, April 3, 1999

FOR DATES: From 7:00 a.m. Friday, April 2, 1999 to 7:00 a.m. Saturday, April 3, 1999.

WEATHER: Overcast. The temperature range was 51oF to 68oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 899.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 37.6 tons per hour while on-line. The contractor is performing related site activities such as feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, ash handling and ash backfilling. Excavation activities were performed in the exclusion zone south of the CZ113 sheet piling. A survey of the excavation bottom and the 2 foot rock bottom was performed. Rock from the screeners was used for backfilling on the 2-foot rock bottom. Continued with the real time perimeter air monitoring (PAM) program and the OSHA method 93 24-hour air sampling event for BNA at the on-site and off-site monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, April 2, 1999

No report available.


Thursday, April 1, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 31, 1999 to 7:00 a.m. Thursday, April 1, 1999.

WEATHER: Clear and mild. The temperature range was 30oF to 73oF.

I. CURRENT SITUATION: OHM continues to work 24-hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 880.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.8 tons per hour while on-line. The contractor is performing related site activities such as feed preparation, debris separation and equipment demobilization.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, wastewater management, feed preparation, debris separation, equipment decontamination, dust control, excavation surveys, ash handling and ash backfilling. Excavation activities were performed in the exclusion zone adjacent to the CZ113 sheet piling and inside of the feed preparation building. The contractor hand picked rock from the screeeners for backfilling on the 2-foot rock bottom. One roll-off containing the skin of the debris separation sprung structure was shipped as non-hazardous waste to a hazardous waste landfill. Continued with the real time perimeter air monitoring (PAM) program. The second and third run of stack testing for metals and particulate was performed, AirRecon demobilized from the site once testing had been completed.

Received bottom ash analytical report (bot-ash-03-27-99-392) for the contents of bin 6. The analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Tony Eicher (FOCUS), Sandra Downs (PADEP), Kurt Davis (TAG).

PREPARED BY: Steve Gibson for Frank Cleaver (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.